OpenStudy (hang254):

The equation of a circle is (x + 1)^2 + (y + 3)^2 = 25. Which graph best represents the circle?

OpenStudy (anonymous):

looks like \((x-h)^2+(y-k)^2=r^2\) identify the center as \((h,k)\) and the radius as \(r\)

OpenStudy (hang254):

@satellite73 how would i apply that to the graphs/

OpenStudy (anonymous):

look for the one with radius \(r\) since \(r^2=25\) you know \(r=5\)

OpenStudy (anonymous):

also you know that the center is \((-1,-3)\)

OpenStudy (anonymous):

therefore pick the circle with center \((-1,3)\) and radius 5

OpenStudy (hang254):

Ok, i will look for that

OpenStudy (hang254):

@satellite73 It would be graph number 4?

OpenStudy (anonymous):

i don't know let me look

OpenStudy (hang254):

alright

OpenStudy (anonymous):

yeah looks good

OpenStudy (hang254):

ok, thanks

OpenStudy (anonymous):

yw